please help me with this problem Thank you in advance
x + Y=12
Y=2X

Answer by tanimachatterjee(60)

here we have 2 equations x+y=12 ...........(1)
and y=2x ..........(2)
now we can replace the value of y of equation (1) by 2x
so we have x+2x =12
or 3x =12
or x = 12/3= 4
we know from equation (2) that y = 2x
therefore value of y = 2*4= 8

ans ..... x = 4 and y = 8